\begin{thebibliography}{}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\providecommand{\url}[1]{\href{#1}{#1}}
\providecommand{\dodoi}[1]{doi:~\href{http://doi.org/#1}{\nolinkurl{#1}}}
\providecommand{\doeprint}[1]{\href{http://ascl.net/#1}{\nolinkurl{http://ascl.net/#1}}}
\providecommand{\doarXiv}[1]{\href{https://arxiv.org/abs/#1}{\nolinkurl{https://arxiv.org/abs/#1}}}

\bibitem[{{Chluba} {et~al.}(2021){Chluba}, {Abitbol}, {Aghanim},
  {Ali-Ha{\"\i}moud}, {Alvarez}, {Basu}, {Bolliet}, {Burigana}, {de Bernardis},
  {Delabrouille}, {Dimastrogiovanni}, {Finelli}, {Fixsen}, {Hart},
  {Hern{\'a}ndez-Monteagudo}, {Hill}, {Kogut}, {Kohri}, {Lesgourgues},
  {Maffei}, {Mather}, {Mukherjee}, {Patil}, {Ravenni}, {Remazeilles}, {Rotti},
  {Rubi{\~n}o-Martin}, {Silk}, {Sunyaev}, \& {Switzer}}]{2021ExA....51.1515C}
{Chluba}, J., {Abitbol}, M.~H., {Aghanim}, N., {et~al.} 2021, Experimental
  Astronomy, 51, 1515, \dodoi{10.1007/s10686-021-09729-5}

\bibitem[{{Fixsen}(2009)}]{2009ApJ...707..916F}
{Fixsen}, D.~J. 2009, \apj, 707, 916, \dodoi{10.1088/0004-637X/707/2/916}

\bibitem[{{Fixsen} {et~al.}(1994){Fixsen}, {Cheng}, {Cottingham}, {Eplee},
  {Isaacman}, {Mather}, {Meyer}, {Noerdlinger}, {Shafer}, {Weiss}, {Wright},
  {Bennett}, {Boggess}, {Kelsall}, {Moseley}, {Silverberg}, {Smoot}, \&
  {Wilkinson}}]{1994ApJ...420..445F}
{Fixsen}, D.~J., {Cheng}, E.~S., {Cottingham}, D.~A., {et~al.} 1994, \apj, 420,
  445, \dodoi{10.1086/173575}

\bibitem[{{Gaia Collaboration} {et~al.}(2021){Gaia Collaboration}, {Klioner},
  {Mignard}, {Lindegren}, {Bastian}, {McMillan}, {Hern{\'a}ndez}, {Hobbs},
  {Ramos-Lerate}, {Biermann}, {Bombrun}, {de Torres}, {Gerlach}, {Geyer},
  {Hilger}, {Lammers}, {Steidelm{\"u}ller}, {Stephenson}, {Brown}, {Vallenari},
  {Prusti}, {de Bruijne}, {Babusiaux}, {Creevey}, {Evans}, {Eyer}, {Hutton},
  {Jansen}, {Jordi}, {Luri}, {Panem}, {Pourbaix}, {Randich}, {Sartoretti},
  {Soubiran}, {Walton}, {Arenou}, {Bailer-Jones}, {Cropper}, {Drimmel}, {Katz},
  {Lattanzi}, {van Leeuwen}, {Bakker}, {Casta{\~n}eda}, {De Angeli},
  {Ducourant}, {Fabricius}, {Fouesneau}, {Fr{\'e}mat}, {Guerra}, {Guerrier},
  {Guiraud}, {Jean-Antoine Piccolo}, {Masana}, {Messineo}, {Mowlavi},
  {Nicolas}, {Nienartowicz}, {Pailler}, {Panuzzo}, {Riclet}, {Roux},
  {Seabroke}, {Sordo}, {Tanga}, {Th{\'e}venin}, {Gracia-Abril}, {Portell},
  {Teyssier}, {Altmann}, {Andrae}, {Bellas-Velidis}, {Benson}, {Berthier},
  {Blomme}, {Brugaletta}, {Burgess}, {Busso}, {Carry}, {Cellino}, {Cheek},
  {Clementini}, {Damerdji}, {Davidson}, {Delchambre}, {Dell'Oro},
  {Fern{\'a}ndez-Hern{\'a}ndez}, {Galluccio}, {Garc{\'\i}a-Lario},
  {Garcia-Reinaldos}, {Gonz{\'a}lez-N{\'u}{\~n}ez}, {Gosset}, {Haigron},
  {Halbwachs}, {Hambly}, {Harrison}, {Hatzidimitriou}, {Heiter}, {Hestroffer},
  {Hodgkin}, {Holl}, {Jan{\ss}en}, {Jevardat de Fombelle}, {Jordan},
  {Krone-Martins}, {Lanzafame}, {L{\"o}ffler}, {Lorca}, {Manteiga}, {Marchal},
  {Marrese}, {Moitinho}, {Mora}, {Muinonen}, {Osborne}, {Pancino}, {Pauwels},
  {Recio-Blanco}, {Richards}, {Riello}, {Rimoldini}, {Robin}, {Roegiers},
  {Rybizki}, {Sarro}, {Siopis}, {Smith}, {Sozzetti}, {Ulla}, {Utrilla}, {van
  Leeuwen}, {van Reeven}, {Abbas}, {Abreu Aramburu}, {Accart}, {Aerts},
  {Aguado}, {Ajaj}, {Altavilla}, {{\'A}lvarez}, {{\'A}lvarez Cid-Fuentes},
  {Alves}, {Anderson}, {Anglada Varela}, {Antoja}, {Audard}, {Baines}, {Baker},
  {Balaguer-N{\'u}{\~n}ez}, {Balbinot}, {Balog}, {Barache}, {Barbato},
  {Barros}, {Barstow}, {Bartolom{\'e}}, {Bassilana}, {Bauchet},
  {Baudesson-Stella}, {Becciani}, {Bellazzini}, {Bernet}, {Bertone}, {Bianchi},
  {Blanco-Cuaresma}, {Boch}, {Bossini}, {Bouquillon}, {Bramante}, {Breedt},
  {Bressan}, {Brouillet}, {Bucciarelli}, {Burlacu}, {Busonero}, {Butkevich},
  {Buzzi}, {Caffau}, {Cancelliere}, {C{\'a}novas}, {Cantat-Gaudin}, {Carballo},
  {Carlucci}, {Carnerero}, {Carrasco}, {Casamiquela}, {Castellani},
  {Castro-Ginard}, {Castro Sampol}, {Chaoul}, {Charlot}, {Chemin}, {Chiavassa},
  {Comoretto}, {Cooper}, {Cornez}, {Cowell}, {Crifo}, {Crosta}, {Crowley},
  {Dafonte}, {Dapergolas}, {David}, {David}, {de Laverny}, {De Luise}, {De
  March}, {De Ridder}, {de Souza}, {de Teodoro}, {del Peloso}, {del Pozo},
  {Delgado}, {Delgado}, {Delisle}, {Di Matteo}, {Diakite}, {Diener},
  {Distefano}, {Dolding}, {Eappachen}, {Enke}, {Esquej}, {Fabre}, {Fabrizio},
  {Faigler}, {Fedorets}, {Fernique}, {Fienga}, {Figueras}, {Fouron},
  {Fragkoudi}, {Fraile}, {Franke}, {Gai}, {Garabato}, {Garcia-Gutierrez},
  {Garc{\'\i}a-Torres}, {Garofalo}, {Gavras}, {Giacobbe}, {Gilmore}, {Girona},
  {Giuffrida}, {Gomez}, {Gonzalez-Santamaria}, {Gonz{\'a}lez-Vidal}, {Granvik},
  {Guti{\'e}rrez-S{\'a}nchez}, {Guy}, {Hauser}, {Haywood}, {Helmi}, {Hidalgo},
  {H{\l}adczuk}, {Holland}, {Huckle}, {Jasniewicz}, {Jonker}, {Juaristi
  Campillo}, {Julbe}, {Karbevska}, {Kervella}, {Khanna}, {Kochoska},
  {Kordopatis}, {Korn}, {Kostrzewa-Rutkowska}, {Kruszy{\'n}ska}, {Lambert},
  {Lanza}, {Lasne}, {Le Campion}, {Le Fustec}, {Lebreton}, {Lebzelter},
  {Leccia}, {Leclerc}, {Lecoeur-Taibi}, {Liao}, {Licata}, {Lindstr{\o}m},
  {Lister}, {Livanou}, {Lobel}, {Madrero Pardo}, {Managau}, {Mann}, {Marchant},
  {Marconi}, {Marcos Santos}, {Marinoni}, {Marocco}, {Marshall}, {Martin Polo},
  {Mart{\'\i}n-Fleitas}, {Masip}, {Massari}, {Mastrobuono-Battisti}, {Mazeh},
  {Messina}, {Michalik}, {Millar}, {Mints}, {Molina}, {Molinaro}, {Moln{\'a}r},
  {Montegriffo}, {Mor}, {Morbidelli}, {Morel}, {Morris}, {Mulone}, {Munoz},
  {Muraveva}, {Murphy}, {Musella}, {Noval}, {Ord{\'e}novic}, {Orr{\`u}},
  {Osinde}, {Pagani}, {Pagano}, {Palaversa}, {Palicio}, {Panahi}, {Pawlak},
  {Pe{\~n}alosa Esteller}, {Penttil{\"a}}, {Piersimoni}, {Pineau}, {Plachy},
  {Plum}, {Poggio}, {Poretti}, {Poujoulet}, {Pr{\v{s}}a}, {Pulone}, {Racero},
  {Ragaini}, {Rainer}, {Raiteri}, {Rambaux}, {Ramos}, {Re Fiorentin}, {Regibo},
  {Reyl{\'e}}, {Ripepi}, {Riva}, {Rixon}, {Robichon}, {Robin}, {Roelens},
  {Rohrbasser}, {Romero-G{\'o}mez}, {Rowell}, {Royer}, {Rybicki}, {Sadowski},
  {Sagrist{\`a} Sell{\'e}s}, {Sahlmann}, {Salgado}, {Salguero}, {Samaras},
  {Sanchez Gimenez}, {Sanna}, {Santove{\~n}a}, {Sarasso}, {Schultheis},
  {Sciacca}, {Segol}, {Segovia}, {S{\'e}gransan}, {Semeux}, {Siddiqui},
  {Siebert}, {Siltala}, {Slezak}, {Smart}, {Solano}, {Solitro}, {Souami},
  {Souchay}, {Spagna}, {Spoto}, {Steele}, {S{\"u}veges}, {Szabados},
  {Szegedi-Elek}, {Taris}, {Tauran}, {Taylor}, {Teixeira}, {Thuillot},
  {Tonello}, {Torra}, {Torra}, {Turon}, {Unger}, {Vaillant}, {van Dillen},
  {Vanel}, {Vecchiato}, {Viala}, {Vicente}, {Voutsinas}, {Weiler}, {Wevers},
  {Wyrzykowski}, {Yoldas}, {Yvard}, {Zhao}, {Zorec}, {Zucker}, {Zurbach}, \&
  {Zwitter}}]{2021A&A...649A...9G}
{Gaia Collaboration}, {Klioner}, S.~A., {Mignard}, F., {et~al.} 2021, \aap,
  649, A9, \dodoi{10.1051/0004-6361/202039734}

\bibitem[{{Kamionkowski} \& {Knox}(2003)}]{2003PhRvD..67f3001K}
{Kamionkowski}, M., \& {Knox}, L. 2003, \prd, 67, 063001,
  \dodoi{10.1103/PhysRevD.67.063001}

\bibitem[{{Kogut} {et~al.}(1993){Kogut}, {Lineweaver}, {Smoot}, {Bennett},
  {Banday}, {Boggess}, {Cheng}, {de Amici}, {Fixsen}, {Hinshaw}, {Jackson},
  {Janssen}, {Keegstra}, {Loewenstein}, {Lubin}, {Mather}, {Tenorio}, {Weiss},
  {Wilkinson}, \& {Wright}}]{1993ApJ...419....1K}
{Kogut}, A., {Lineweaver}, C., {Smoot}, G.~F., {et~al.} 1993, \apj, 419, 1,
  \dodoi{10.1086/173453}

\bibitem[{{Lange} \& {Page}(2007)}]{2007ApJ...671.1075L}
{Lange}, S., \& {Page}, L. 2007, \apj, 671, 1075, \dodoi{10.1086/523097}

\bibitem[{{Loeb} \& {Narayan}(2008)}]{2008MNRAS.386.2221L}
{Loeb}, A., \& {Narayan}, R. 2008, \mnras, 386, 2221,
  \dodoi{10.1111/j.1365-2966.2008.13187.x}

\bibitem[{{Moss} {et~al.}(2008){Moss}, {Zibin}, \&
  {Scott}}]{2008PhRvD..77d3505M}
{Moss}, A., {Zibin}, J.~P., \& {Scott}, D. 2008, \prd, 77, 043505,
  \dodoi{10.1103/PhysRevD.77.043505}

\bibitem[{{Notari} \& {Quartin}(2015)}]{2015JCAP...06..047N}
{Notari}, A., \& {Quartin}, M. 2015, \jcap, 2015, 047,
  \dodoi{10.1088/1475-7516/2015/06/047}

\bibitem[{{Zibin} {et~al.}(2007){Zibin}, {Moss}, \&
  {Scott}}]{2007PhRvD..76l3010Z}
{Zibin}, J.~P., {Moss}, A., \& {Scott}, D. 2007, \prd, 76, 123010,
  \dodoi{10.1103/PhysRevD.76.123010}

\end{thebibliography}

\documentclass[modern]{aastex61}

\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{enumitem}

\newcommand\mybar{\kern1pt\rule[-\dp\strutbox]{.8pt}{\baselineskip}\kern1pt}

\setlist[itemize]{noitemsep, topsep=0pt, leftmargin=*}

\shorttitle{Evolution of the CMB Quadrupole}
\shortauthors{Loeb}

\begin{document}

\title{Time Evolution of the CMB Quadrupole}

\author{Abraham Loeb}
\affiliation{Astronomy Department, Harvard University, 60 Garden
  St., Cambridge, MA 02138, USA}

\begin{abstract}
I show that the quadrupole of the Cosmic Microwave Background (CMB)
evolves more rapidly than previously expected, as a result of the
acceleration of the Sun towards the Galactic center. The acceleration,
measured most recently by {\it Gaia} EDR3, implies a fractional change
in the quadrupole of $\sim 10^{-9}$ per year, an order of magnitude
larger than expected from the evolution in the last scattering surface
of the CMB.

\end{abstract}

*\section*{}
\section*{Introduction}
Within a timescale of order the age of the Universe, $\sim 10^{10}$~yr,
the large-scale anisotropies of the Cosmic Microwave Background (CMB)
are expected to change as a result of the change in the location of
the last scaterring surface \citep{2007ApJ...671.1075L,2007PhRvD..76l3010Z,2008PhRvD..77d3505M}. Naively, one
would expect anisotropies on the largest scales - such as the
quadrupole moment, to change most slowly. However, here we show this
not to be true as a result of the acceleration of the Sun towards the
center of the Milky Way. This acceleration was measured most recently
by {\it Gaia} EDR3 to be, ${\dot v}=2.32(\pm0.16)\times 10^{-8}~{\rm
  cm~s^{-2}}$, towards the Galactic center
\citep{2021A&A...649A...9G}.

\section*{Calculation}

The velocity of the Sun relative to the cosmic frame of reference,
${\vec {\bf v}}$, results in a CMB dipole moment, which is routinely
removed from CMB anistropy maps. But to second-order in ${\vec {\bf
    \beta}}\equiv ({\vec {\bf v}}/c)$, the motion also leads to a
quadrupole anisotropy with an angular dependence of $(\cos^2\theta
-1/3)$, where $\theta$ is the angle between the velocity vector ${\vec
  {\bf v}}$ and the photon direction.

The fractional change in the CMB intensity as a result of the
kinematic quadrupole depends on photon frequency
\citep{2003PhRvD..67f3001K},
\begin{equation}
\left({\Delta I_\nu\over I_\nu}\right)_Q=F(x) \beta^2 ,
\label{one}
\end{equation}
with $F(x)=[xe^x/(e^x-1)][(x/2)\coth(x/2)]$. Here, $x=(h\nu/kT)$ with
$I_\nu=[2(kT)^3/(hc)^2] x^3/(e^x-1)$ being the CMB blackbody intensity
at a frequency $\nu$ and the mean CMB temperature $T=2.725$K
\citep{2009ApJ...707..916F}. At the low frequencies of the
Rayleigh-Jeans regime where $x\ll1$, we get $F(x)\approx 1$ and
$(\Delta I_\nu/I_\nu)=(\Delta T/T)$.

\section*{Conclusions}

The time derivative of equation (\ref{one}) yields a kinematic
variation in the quadrupole, which for $x\ll1$ is given by,
\begin{equation}
\left({{\dot T}\over T}\right)_Q= 2{\vec {\bf \beta}}\cdot{\dot {\vec {\bf \beta}}} ,
\label{two}
\end{equation}
where from the {\it Gaia} measurement, ${\dot \beta}=2.44\times
10^{-11}~{\rm yr^{-1}}$, and from the CMB dipole $\beta \approx
1.23\times 10^{-3}$
\citep{1993ApJ...419....1K,1994ApJ...420..445F}. The dot-product of
the acceleration vector towards the Galactic center and the CMB dipole
velocity vector yields a geometric projection factor of $0.065$
\citep{2008MNRAS.386.2221L}.

Given the quadrupole moment measured by the Planck satellite, $Q\equiv
(\Delta T/T)_Q\sim 4.5\times 10^{-6}$ \citep{2015JCAP...06..047N}, the
fractional time derivative of the CMB quadrupole, as a result of the
acceleration of the Sun towards the Galactic center, is given by,
\begin{equation}
\left({{\dot Q}\over Q}\right) \approx 10^{-9}~{\rm yr^{-1}} .
\label{three}
\end{equation}
This rate of evolution is over an order of magnitude larger than
expected from the change in the last scattering surface of the CMB
\citep{2007ApJ...671.1075L,2007PhRvD..76l3010Z,2008PhRvD..77d3505M}.

The expected time evolution of the CMB quadrupole spectrum, $F(x)$, in
equation (\ref{one}), could potentially be measured - especially for
$x\gg 1$ where $F(x)\gg 1$, with future CMB spectral distortion
experiments \citep{2021ExA....51.1515C}.

\bigskip
\bigskip
\section*{Acknowledgements}

This work was supported in part by Harvard's {\it Black Hole
  Initiative}, which is funded by grants from JFT and GBMF.

\bibliographystyle{aasjournal}
\bibliography{Q}
\label{lastpage}
\end{document}